# A Systematic Literature Review of Spatio-Temporal Graph Neural Network Models for Time Series Forecasting and Classification

### v2

@@ 5.6.3 Types of models
More than half of the proposed spatio-temporal GNN models (19 out of 35 papers) fall under the category of pure convolutional GNNs, followed by 10 attentional GNNs, 2 hybrid models, and one recurrent model (Ref. [art_470]). One paper (Ref. [art_57]) uses both convolutional and attentional GNNs, one paper (Ref. [art_475]) uses a variant of the convolutional GNN, and the description of another one (Ref. [art_155]) does not clarify which class of the taxonomy it belongs to. Almost all the models have a graph learning module, so that the graph structure is rarely defined a priori by the researchers. Three papers [art_55, art_57, art_154] use the visibility graph algorithm proposed by Lacasa et al. in Ref. [visibility_graph] for converting the time series into a graph.
@@
Convolution-graph attention network (ConvGAT) [art_153] https://github.com/CUG-FDGroup
Graph Network with Causal Connectivity and Temporal Convolutional Neural Network feature extractor [art_155] https://github.com/mylonasc/gnn-tcnn
Graph spatiotemporal process and anomaly scorer (GST-Pro) [art_477] https://github.com/huankoh/GST-Pro
Causality Enhanced Global-Local Graph Neural Network (CEGLo-GNN) [art_482] https://github.com/YueYueXia/CEGLo-GNN
Graph Deviation Network+ (GDN+) [art_483] https://github.com/KatieBuc/gnnad/releases
@@ 5.6.4 Benchmark models
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_483], [art_487]
Non-GNN machine learning methods Attention-based long-short term memory network [art_153], [art_157], [art_488]
Feature-attention based bidirectional GRU and CNN model (AGCNN) [AGCNN] [art_153], [art_154], [art_157], [art_478], [art_488]
@@
Anomalous Rhythm Detection using Adversarially Generated Time Series (BeatGAN) [BeatGAN] [art_481], [art_484]
Bidirectional long-short term memory network (BiLSTM) [BiLSTM] [art_152], [art_153], [art_154], [art_157], [art_470], [art_488]
Convolutional neural network (CNN) [art_9], [art_56], [art_152], [art_153], [art_154], [art_156], [art_157], [art_461], [art_469], [art_471], [art_486], [art_487], [art_488]
@@
Convolutional neural network with long-short term memory network (CNN-LSTM) [art_154], [art_470]
Crossformer [Crossformer] [art_152], [art_482]
Deep autoencoding gaussian model (DAGMM) [DAGMM] [art_481], [art_484]
Deep belief network (DBN) [DBN] [art_157], [art_488]
Decision tree (DT) [art_9], [art_56], [art_487]
Feed-forward neural network (FNN) [art_55], [art_155], [art_469], [art_478]
@@
Genetic algorithm approach on restricted Boltzmann-based long-short term memory network (GA+RBM+LSTM) [GA+RBM+LSTM] [art_157], [art_488]
Gradient boosting (GB) [art_56], [art_154], [art_487]
Gated recurrent unit (GRU) [art_53], [art_56], [art_156], [art_469], [art_487]
@@
Integrated multi-head dual sparse self-attention network (IMDSSN) [IMDSSN] [art_154], [art_478], [art_488]
Isolation forest [art_10], [art_481]
$k$-nearest neighbors (KNN) [art_56], [art_475], [art_487]
Local outlier factor (LOF) [art_475], [art_481]
Logistic regression (LR) [art_9], [art_56]
Long-short term memory network (LSTM) [art_55], [art_56], [art_154], [art_157], [art_158], [art_461], [art_470], [art_487], [art_488]
@@
Long-short term memory-based variational auto-encoder (LSTM-VAE) [art_10], [art_477], [art_479], [art_484], [art_489]
OmniAnomaly [OmniAnomaly] [art_477], [art_479], [art_481], [art_484], [art_489]
Residual network (ResNet) [ResNet] [art_471], [art_491]
Random forest (RF) [art_56], [art_154], [art_478], [art_487]
Support vector machine (SVM) [art_9], [art_56], [art_471], [art_478], [art_486], [art_487]
Deep transformer networks for anomaly detection (TranAD) [TranAD] [art_484], [art_489]
Transformer [attention] [art_152], [art_154], [art_470], [art_478], [art_479], [art_488], [art_489]
@@
Unsupervised anomaly detection on multivariate time series (USAD) [5] [art_479], [art_484], [art_489]
GNN methods Adaptive graph convolutional recurrent network (AGCRN) [AGCRN] [art_58], [art_482]
Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_480], [art_482]
Fused sparse autoencoder and graph network (FuSAGNet) [FuSAGNet] [art_10], [art_479]
Graph attention network (GAT) [GAT] [art_54], [art_56], [art_192], [art_470], [art_487]
Graph convolutional network (GCN) [art_54], [art_55], [art_156], [art_192], [art_461], [art_474], [art_486]
Graph deviation network (GDN) [GDN] [art_10], [art_477], [art_479], [art_481], [art_484]
Gated graph convolutional network (GGCN) [GGCN] [art_153], [art_478]
Graph WaveNet (GWN) [GWN] [art_58], [art_480], [art_482]
Hierarchical attention graph convolutional network (HAGCN) [HAGCN] [art_152], [art_153], [art_478]
@@
Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_470], [art_483]
Spatio-temporal fusion attention (STFA) [art_157] [art_153], [art_154], [art_478], [art_488]
Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_58], [art_152], [art_470], [art_478], [art_480]
@@
Among the GNN benchmarks, some of them have been specifically developed for remaining useful life prediction or predictive monitoring tasks, such as the hierarchical attention graph convolutional network (HAGCN) [HAGCN] and the spatio-temporal fusion attention (STFA) [art_157]. In the HAGCN model, proposed by Li et al. in 2021, the spatial dependencies are modeled by a hierarchical graph representation layer, and a bidirectional long-short term memory network is used for modeling temporal dependencies of sensor measurements. As for the STFA model, proposed by Kong et al. in 2022, it integrates some a priori knowledge about the equipment’s structure with a spatio-temporal deep learning architecture which employs LSTM cells and an attention mechanism. Neither of the two original papers includes a link to the source code.
@@ 5.6.5 Results
The only results that can be compared across papers are those for the C-MAPSS dataset. Tab. LABEL:t:risultati_c-mapss shows the accuracy of the different models, expressed in terms of RMSE and score function.
@@
In general, the best results on the different sub-datasets are achieved by the convolution-graph attention network (ConvGAT) [art_153] and the the dual-stream spatio-temporal fusion network (DS-STFN) [art_478] models. The ConvGAT model, proposed by Chen et al. in 2023, applies independent convolutional kernels to each feature of the input time series, and passes the resulting values to a graph attention network, with graph structure defined using cosine similarity. Finally, a fully connected layer outputs the RUL prediction. The DS-STFN model, proposed by Zhang et al. in 2024, organizes the data into grid-structured and graph-structured formats, with the graph structure defined by cosine similarity. A temporal convolutional network with self-attention extracts temporal features from the grid data, and, in parallel, a convolutional graph transformer module captures spatial features from the graph data. The spatial and temporal features are then fused to predict the RUL.
@@ 5.7.1 Overview
A significant number of papers (44 out of 281) have been categorized under the “Generic” thematic group. The majority of these works commit themselves to address a multivariate time series forecasting problem not necessarily limited to a specific field. This “Generic” set includes one paper from 2021, 4 from 2022, 19 from 2023, and 20 from 2024, highlighting the recent interest in applying GNNs in cross-disciplinary contexts. These papers are scattered across various journals, with the exception of Expert Systems with Applications by Elsevier, IEEE Transactions on Knowledge and Data Engineering by IEEE, and Information Sciences by Elsevier, which published 4 and 3 of the selected papers, respectively. The majority of authors are affiliated with institutions in China.
@@ 5.7.2 Datasets
Beijing Traffic [art_84] https://github.com/BuaaPercy/Traffic-DataSet-of-Beijing-Road
CCMP wind data [art_73] http://data.remss.com
CDC US States [art_401] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html
CI earthquakes [art_160] https://doi.org/10.5194/adgeo-43-31-2016 (source paper)
CityBike NYC [art_464] https://www.citibikenyc.com/system-data
Chickenpox cases [art_160] https://doi.org/10.48550/arXiv.2102.08100 (source paper)
C-MAPSS [art_403] https://ntrs.nasa.gov/api/citations/20070034949/downloads/20070034949.pdf
CORAv2.0 [art_396] https://mds.nmdis.org.cn/pages/home.html
CW earthquakes [art_89] https://doi.org/10.5194/adgeo-43-31-2016 (source paper)
ECG5000 [art_397] https://www.cs.ucr.edu/ẽamonn/time_series_data/
@@
Electricity consumption [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_88], [art_90], [art_384], [art_405] https://github.com/laiguokun/multivariate-time-series-data
Electricity Consuming Load (ECL) [art_82], [art_85] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014
Electricity transformer Temperature (ETT) [art_76], [art_82], [art_85], [art_393], [art_417] https://github.com/zhouhaoyi/ETDataset
Energy [art_77], [art_388] http://dx.doi.org/10.1016/j.enbuild.2017.01.083 (source paper)
Enron [art_202] https://doi.org/10.1073/pnas.1800683115 (source paper)
Eu-Core [art_202] https://doi.org/10.1145/3018661.3018731 (source paper)
Exchange-rate [art_73], [art_74], [art_76], [art_77], [art_78], [art_79], [art_80], [art_81], [art_85], [art_87], [art_89], [art_90], [art_384], [art_393], [art_405] https://github.com/laiguokun/multivariate-time-series-data
Facebook [art_202] http://networkrepository.com/socfb
GEFCOM 2012 [art_89] https://doi.org/10.1016/j.ijforecast.2013.07.001 (source paper)
Hypertext [art_202] http://networkrepository.com/ia-infect-hyper (source paper)
IDWR Japan Prefectures [art_401] https://www.niid.go.jp/niid/en/idwr-e.html
ILI [art_76], [art_85], [art_89] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html
ISRUC-SLEEP Dataset [art_403] https://sleeptight.isr.uc.pt/ISRUC_Sleep/
Jena Weather [art_76] https://www.bgc-jena.mpg.de/wetter/
KnowAir [art_400] https://github.com/shuowang-ai/PM2.5-GNN
LA Traffic [art_404] https://ladot.lacity.gov/residents/traffic-counts
Loan [art_202] https://ai.ppdai.com/mirror/showCompetitionRisk
Los-loop [art_383] https://github.com/lehaifeng/T-GCN/tree/master/data
Large-Scale Traffic and Weather Events (LSTW) [art_404] https://smoosavi.org/datasets/lstw
METR-LA [art_72], [art_75], [art_79], [art_80], [art_86], [art_160], [art_387], [art_397], [art_401], [art_417], [art_464] https://paperswithcode.com/dataset/metr-la
Metro [art_400] https://lbs.amap.com/
MIMIC-III Heart Failure [art_71] https://doi.org/10.1038/sdata.2016.35 (source paper)
MIMIC-III Infection [art_71] https://doi.org/10.1038/sdata.2016.35 (source paper)
MOOC network [art_70] http://snap.stanford.edu/jodie/mooc.csv
Nasdaq [art_77], [art_80] https://doi.org/10.48550/arXiv.1704.02971 (source paper)
NOAA Weather [art_82], [art_393] https://www.ncei.noaa.gov/data/local-climatological-data/
NYC-Taxi [art_398] https://www1.nyc.gov/site/tlc/about/tlc-trip-record-data.page
PeMS03 [art_74] https://www.kaggle.com/datasets/elmahy/pems-dataset
PeMS04 (PeMSD4) [art_73], [art_74], [art_78], [art_79], [art_88], [art_384], [art_396], [art_397], [art_400], [art_405] https://paperswithcode.com/dataset/pems04
PeMS08 (PeMSD8) [art_73], [art_74], [art_78], [art_79], [art_88], [art_383], [art_384], [art_396], [art_397], [art_400], [art_405] https://paperswithcode.com/dataset/pems08
PEMS-BAY [art_72], [art_75], [art_160], [art_401] https://paperswithcode.com/dataset/pems-bay
PhysioNet [art_71] https://archive.physionet.org/challenge/2012/papers/
Reddit network [art_70] http://snap.stanford.edu/jodie/reddit.csv
Seattle-loop [art_401] https://github.com/zhiyongc/Seattle-Loop-Data
Sleep-EDF-20 [art_395] https://gist.github.com/emadeldeen24/a22691e36759934e53984289a94cb09b
Solar-Energy [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_87], [art_88], [art_90], [art_384], [art_387], [art_405] https://github.com/laiguokun/multivariate-time-series-data
Secure Water Treatment (SWaT) [art_387] https://itrust.sutd.edu.sg/itrust-labs_datasets/
Stocktwits [art_398] https://stocktwits.com/
TAIEX [art_84] https://finance.yahoo.com/
TaoBao [art_202] https://tianchi.aliyun.com/competition/entrance/
Taiwan PM2.5 [art_464] https://taqm.epa.gov.tw/taqm/en/default.aspx
T-drive [art_383] https://www.kaggle.com/datasets/arashnic/tdriver
Tiingo [art_398] https://www.tiingo.com/
Traffic [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_85], [art_87], [art_90], [art_388] https://github.com/laiguokun/multivariate-time-series-data
UCI Human Activity Recognition [art_403] https://archive.ics.uci.edu/dataset/240/human+activity+recognition+using+smartphones
University of East Anglia (UEA) datasets [art_385], [art_386], [art_389], [art_407] https://timeseriesclassification.com/
US stock market price [art_89] https://doi.org/10.48550/arXiv.1810.09936 (source paper)
Water Distribution (WaDi) [art_387] https://itrust.sutd.edu.sg/itrust-labs_datasets/
WESAD dataset [art_404] https://ubi29.informatik.uni-siegen.de/usi/data_wesad.html
Wikipedia network [art_70] http://snap.stanford.edu/jodie/wikipedia.csv
Wind-Speed [art_76] https://www.kaggle.com/datasets/fedesoriano/wind-speed-prediction-dataset
@@
Wind Generation [art_405] https://www.kaggle.com/datasets/sohier/30-years-of-european-wind-generation
@@
WISDM [art_395] https://archive.ics.uci.edu/dataset/507/wisdm+smartphone+and+smartwatch+activity+and+biometrics+dataset)
@@
The University of East Anglia (UEA) dataset includes 30 multivariate time series from many real-world applications, including heart records, electroencephalogram signals, motion data, audio spectra, written characters. Since a large portion of the dataset contains data not related to the focus of this review, the details of the results and models used will not be discussed here. Interested readers are referred to Refs. [art_385, art_386, art_389, art_407].

### v4

@@ 5.6.3 Types of models
Almost half of the proposed spatio-temporal GNN models (19 out of 40 papers) fall under the category of pure convolutional GNNs, followed by 15 attentional GNNs, 2 hybrid models, and one recurrent model (Ref. [86]). One paper (Ref. [296]) uses both convolutional and attentional GNNs, one paper (Ref. [96]) uses a variant of the convolutional GNN, and the description of another one (Ref. [291]) does not clarify which class of the taxonomy it belongs to. Almost all the models have a graph learning module, so that the graph structure is rarely defined a priori by the researchers. Three papers [207, 296, 382] use the visibility graph algorithm proposed by Lacasa et al. in Ref. [201] for converting the time series into a graph.
@@
Causality enhanced global-local graph neural network (CEGLo-GNN) [363] https://github.com/YueYueXia/CEGLo-GNN
Dual time-oriented graph attention networks (DuoGAT) [204] https://github.com/ByeongtaePark/DuoGAT
Graph deviation network+ (GDN+) [21] https://github.com/KatieBuc/gnnad/releases
Graph network with causal connectivity and temporal convolutional neural network feature extractor [291] https://github.com/mylonasc/gnn-tcnn
Spectral multi-view causal graph fusion based anomaly detection (SMV-CGAD) [283] https://github.com/arunvignesh28/SMV-CGAD
Spatio-temporal graph neural network empowered long-term trend analysis system (ST-GTrend) [79] https://github.com/Yangxin666/ST-GTrend
Graph attention networks with topological analysis (TopoGDN) [272] https://github.com/ljj-cyber/TopoGDN
@@ 5.6.4 Benchmark models
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [21], [476]
Non-GNN machine learning methods Auto-encoder (AE) [512], [128], [204], [139]
Feature-attention based bidirectional GRU and CNN model (AGCNN) [526] [46], [382], [195], [488], [194]
Attention-based long-short term memory network [46], [195], [194]
@@
Anomalous rhythm detection using adversarially generated time series (BeatGAN) [518] [114], [512]
Bidirectional long-short term memory network (BiLSTM) [109] [400], [46], [382], [195], [86], [194]
Convolutional neural network (CNN) [221], [400], [46], [382], [441], [195], [31], [196], [379], [476], [194]
@@
Convolutional neural network with long-short term memory network (CNN-LSTM) [382], [86]
Crossformer [503] [400], [363]
Deep autoencoding gaussian model (DAGMM) [529] [114], [512], [128], [239], [204]
Deep belief network (DBN) [474] [195], [194]
Decision tree (DT) [221], [476]
Feed-forward neural network (FNN) [207], [291], [196], [488]
@@
Genetic algorithm approach on restricted Boltzmann-based long-short term memory network (GA+RBM+LSTM) [258] [195], [194]
Gradient boosting (GB) [382], [476]
Gated recurrent unit (GRU) [374], [441], [196], [476]
@@
Integrated multi-head dual sparse self-attention network (IMDSSN) [478] [382], [488], [194]
Isolation forest [226], [114], [283]
$k$-nearest neighbors (KNN) [96], [476], [128], [204]
Local outlier factor (LOF) [19] [96], [114]
Long-short term memory network (LSTM) [207], [382], [195], [408], [31], [86], [476], [194]
@@
Long-short term memory-based variational auto-encoder (LSTM-VAE) [226], [515], [249], [512], [136], [128], [272], [139]
Multivariate anomaly detection with GAN (MAD-GAN) [209] [226], [128], [272], [204]
OmniAnomaly [358] [515], [249], [114], [512], [136], [272]
Random forest (RF) [382], [488], [476]
Support vector machine (SVM) [221], [128], [488], [379], [476], [283], [204], [139]
Deep transformer networks for anomaly detection (TranAD) [372] [512], [136], [272], [239]
Transformer [375] [400], [382], [86], [488], [249], [194], [136]
Unsupervised anomaly detection on multivariate time series (USAD) [4] [249], [512], [136], [128], [204], [139]
GNN methods Adaptive graph convolutional recurrent network (AGCRN) [8] [497], [363]
Diffusion convolutional recurrent neural network (DCRNN) [237] [224], [363]
Fused sparse autoencoder and graph network (FuSAGNet) [128] [226], [249], [239]
Graph attention network (GAT) [376] [464], [206], [86], [476]
Graph convolutional network (GCN) [464], [207], [441], [206], [31], [389], [379]
Graph deviation network (GDN) [66] [226], [515], [249], [114], [512], [128], [272], [239], [204], [139]
Gated graph convolutional network (GGCN) [386] [46], [488]
Graph WaveNet (GWN) [425] [497], [224], [363]
Hierarchical attention graph convolutional network (HAGCN) [230] [400], [46], [488]
@@
Multivariate time-series anomaly detection via graph attention network (MTAD-GAT) [507] [249], [128], [239], [204]
@@
Multivariate time series forecasting with graph neural network (MTGNN) [424] [86], [21]
Spatio-temporal fusion attention (STFA) [195] [46], [382], [488], [194]
Spatio-temporal graph convolutional network (STGCN) [457] [497], [400], [86], [488], [224]
@@
Among the GNN benchmarks, some of them have been specifically developed for remaining useful life prediction or predictive monitoring tasks, such as the hierarchical attention graph convolutional network (HAGCN) [230], the spatio-temporal fusion attention (STFA) [195], and the graph deviation network (GCN) [66]. In the HAGCN model, proposed by Li et al. in 2021, the spatial dependencies are modeled by a hierarchical graph representation layer, and a bidirectional long-short term memory network is used for modeling temporal dependencies of sensor measurements. As for the STFA model, proposed by Kong et al. in 2022, it integrates some a priori knowledge about the equipment’s structure with a spatio-temporal deep learning architecture which employs LSTM cells and an attention mechanism. Finally, the GDN model, proposed by Deng et al. in 2021, learns a graph structure that represents the relationships between sensors, and aims to detect deviations from these patterns by using an attentional GNN. The source code for the GDN model is available at https://github.com/d-ailin/GDN.
@@ 5.6.5 Results
The only results that can be compared across many papers are those for the C-MAPSS dataset. Tab. LABEL:t:risultati_c-mapss shows the accuracy of the different models, expressed in terms of RMSE and score function.
@@
In general, the best results on the different sub-datasets are achieved by the convolution-graph attention network (ConvGAT) [46] and the the dual-stream spatio-temporal fusion network (DS-STFN) [488] models. The ConvGAT model, proposed by Chen et al. in 2023, applies independent convolutional kernels to each feature of the input time series, and passes the resulting values to a graph attention network, with graph structure defined using cosine similarity. Finally, a fully connected layer outputs the RUL prediction. The DS-STFN model, proposed by Zhang et al. in 2024, organizes the data into grid-structured and graph-structured formats, with the graph structure defined by cosine similarity. A temporal convolutional network with self-attention extracts temporal features from the grid data, and, in parallel, a convolutional graph transformer module captures spatial features from the graph data. The spatial and temporal features are then fused to predict the RUL.
Other pairs of papers with comparable results are Refs. [128] and [204] for SWaT and WaDi datasets and Refs. [114] and [272] for the MSL dataset. However, due to space limitations, tables comparing just two studies are not included here.
@@ 5.7 Generic
For instance, some subsets of papers within the “Generic” group focus on forecasting in presence of missing data (e.g., Refs. [458, 286]), or of irregularly sampled time series (e.g., Refs. [436, 499, 491, 493, 300]). Notably, the majority of these studies have been published in conferences rather than in journals.
Moreover, some papers in this section also tackle some theoretical issues. For example, Ref. [271] proves a theorem that provides a theoretical bound on the approximation error of the module that learns the dynamic variable dependence. Another example can be found in Ref. [505], where the authors theoretically establish that the generalization error bound of their model is related to various components of the GAT model, such as the number of attention heads, the number of neighboring nodes, the Frobenius norm of the weight matrices, and the norm of the input features.
@@ 5.7.1 Overview
A significant number of papers (77 out of 366) have been categorized under the “Generic” thematic group. The majority of these works commit themselves to address a multivariate time series forecasting problem not necessarily limited to a specific field. This “Generic” set includes 4 papers from 2002, 7 from 2021, 11 from 2022, 28 from 2023, and 27 from 2024, highlighting the recent interest in applying GNNs in cross-disciplinary contexts. These papers are scattered across various journals, with the exception of Expert Systems with Applications by Elsevier, IEEE Transactions on Knowledge and Data Engineering by IEEE, and Information Sciences by Elsevier, which published 4 and 3 of the selected papers, respectively. The three conference venues with the highest number of papers are NeurIPS with 6 papers, CIKM with 5 papers, and KDD with 4 papers.
@@ 5.7.2 Datasets
Beijing Traffic [47] https://github.com/BuaaPercy/Traffic-DataSet-of-Beijing-Road
Bytom [50], [51], [48] https://etherscan.io/
CCMP wind data [241] http://data.remss.com
CDC US States [360] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html
CER-E [59] https://www.ucd.ie/issda/data/commissionforenergyregulationcer/
Chickenpox [15], [327] https://www.kaggle.com/datasets/die9origephit/chickenpox-cases-hungary
China AQI [458] https://quotsoft.net/air/
CI earthquakes [15] https://doi.org/10.5194/adgeo-43-31-2016 (source paper)
C-MAPSS [401] https://ntrs.nasa.gov/api/citations/20070034949/downloads/20070034949.pdf
CORAv2.0 [137] https://mds.nmdis.org.cn/pages/home.html
COVID-19 CA [51], [48], [451], [49] https://www.dropbox.com/sh/n0ajd5l0tdeyb80/AABGn-ejfV1YtRwjf_L0AOsNa?dl=0
COVID-19 England [327] https://pytorch-geometric-temporal.readthedocs.io/en/latest/notes/introduction.html#benchmark-datasets
COVID-19 TX [51], [48] https://www.dropbox.com/sh/n0ajd5l0tdeyb80/AABGn-ejfV1YtRwjf_L0AOsNa?dl=0
CSSE COVID-19 [432], [451] https://github.com/CSSEGISandData/COVID-19
CW earthquakes [57] https://doi.org/10.5194/adgeo-43-31-2016 (source paper)
Decentraland [50], [51] https://etherscan.io/
ECG5000 [152], [432], [26], [451] https://www.cs.ucr.edu/~eamonn/time_series_data/
@@
Electricity consumption [241], [99], [177], [243], [118], [40], [287], [355], [473], [138], [240], [83], [448], [279], [424], [61], [26], [271], [153], [451], [268] https://github.com/laiguokun/multivariate-time-series-data
Electricity consuming load (ECL) [395], [186], [356] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014
Electricity transformer temperature (ETT) [504], [395], [186], [231], [443], [153], [356] https://github.com/zhouhaoyi/ETDataset
Energy [76], [125] http://dx.doi.org/10.1016/j.enbuild.2017.01.083 (source paper)
ENG-RAD [286] https://open-meteo.com/
Enron [155] https://doi.org/10.1073/pnas.1800683115 (source paper)
Eu-Core [155] https://doi.org/10.1145/3018661.3018731 (source paper)
@@
European wind generation [448] https://www.kaggle.com/datasets/sohier/30-years-of-european-wind-generation
Exchange-rate [241], [99], [504], [76], [243], [118], [40], [287], [186], [397], [57], [138], [231], [240], [448], [424], [61], [271], [153], [268] https://github.com/laiguokun/multivariate-time-series-data
Facebook [155] http://networkrepository.com/socfb
GEFCOM 2012 [57] https://doi.org/10.1016/j.ijforecast.2013.07.001 (source paper)
Golem [51] https://www.dropbox.com/sh/n0ajd5l0tdeyb80/AABGn-ejfV1YtRwjf_L0AOsNa?dl=0
GraphMSO [286] https://github.com/marshka/hdtts/tree/main/config/dataset
Hypertext [155] http://networkrepository.com/ia-infect-hyper (source paper)
HZMetro [279] https://github.com/HCPLab-SYSU/PVCGN/tree/master/data
IDWR Japan prefectures [360] https://www.niid.go.jp/niid/en/idwr-e.html
ILI [504], [186], [57] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html
ISRUC-SLEEP [401] https://sleeptight.isr.uc.pt/ISRUC_Sleep/
Jena weather [504], [153] https://www.bgc-jena.mpg.de/wetter/
KnowAir [390] https://github.com/shuowang-ai/PM2.5-GNN
LA Traffic [337] https://ladot.lacity.gov/residents/traffic-counts
Loan [155] https://ai.ppdai.com/mirror/showCompetitionRisk
Los-loop [244] https://github.com/lehaifeng/T-GCN/tree/master/data
Large-scale traffic and weather events (LSTW) [337] https://smoosavi.org/datasets/lstw
METR-LA [60], [177], [118], [40], [52], [15], [396], [152], [360], [443], [338], [458], [83], [277], [424], [26], [451], [59], [300] https://www.kaggle.com/datasets/annnnguyen/metr-la-dataset
Metro [390] https://lbs.amap.com/
MIMIC [129], [436], [491] https://doi.org/10.1038/sdata.2016.35 (source paper)
Montevideo buses [327] https://pytorch-geometric-temporal.readthedocs.io/en/latest/notes/introduction.html#benchmark-datasets
MOOC network [232] http://snap.stanford.edu/jodie/mooc.csv
MTM-1 hand motions [327] https://pytorch-geometric-temporal.readthedocs.io/en/latest/notes/introduction.html#benchmark-datasets
Nasdaq [76], [40] https://dl.acm.org/doi/10.5555/3172077.3172254 (source paper)
Neonatal EEG [83] https://zenodo.org/records/2547147
NOAA weather [395], [231] https://www.ncei.noaa.gov/data/local-climatological-data/
NREL [432], [286], [451] https://www.nrel.gov/
NYC-Bike [448], [279], [277], [49] https://citibikenyc.com/system-data
NYC-Taxi [453], [450], [448], [279], [49] https://www1.nyc.gov/site/tlc/about/tlc-trip-record-data.page
PAMAP2 [289], [499], [456] https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
Pedal Me deliveries [327] https://pytorch-geometric-temporal.readthedocs.io/en/latest/notes/introduction.html#benchmark-datasets
PeMS03 (PeMSD3) [99], [51], [26], [143] https://www.kaggle.com/datasets/elmahy/pems-dataset
PeMS04 (PeMSD4) [241], [99], [243], [118], [473], [137], [152], [390], [240], [458], [50], [51], [26], [48], [369], [368] https://www.kaggle.com/datasets/elmahy/pems-dataset
PeMS07 (PeMSD7) [26], [271], [369], [368] https://www.kaggle.com/datasets/elmahy/pems-dataset
PeMS08 (PeMSD8) [241], [99], [243], [118], [473], [244], [137], [152], [390], [240], [458], [50], [51], [26], [48], [409], [369], [368] https://www.kaggle.com/datasets/elmahy/pems-dataset
PEMS-BAY [60], [177], [15], [360], [338], [458], [83], [424], [26], [271], [59], [143], [300] https://zenodo.org/records/4263971
PhysioNet [129], [436], [499], [491] https://archive.physionet.org/challenge/2012/papers/
Reddit network [232] http://snap.stanford.edu/jodie/reddit.csv
Seattle-loop [360] https://github.com/zhiyongc/Seattle-Loop-Data
Sleep-EDF-20 [100] https://gist.github.com/emadeldeen24/a22691e36759934e53984289a94cb09b
Solar-Energy [241], [99], [177], [243], [118], [40], [287], [355], [397], [473], [138], [396], [240], [83], [448], [424], [61], [26], [271], [268] https://github.com/laiguokun/multivariate-time-series-data
Secure water treatment (SWaT) [396] https://itrust.sutd.edu.sg/itrust-labs_datasets/
SHMetro [279] https://github.com/HCPLab-SYSU/PVCGN/tree/master/data
Stock market data [505] https://www.kaggle.com/datasets/paultimothymooney/stock-market-data
Stocktwits [453] https://stocktwits.com/
Synthetic periodic data synthetic [300] https://github.com/joeloskarsson/tgnn4i
TAIEX [47] https://finance.yahoo.com/
Taiwan PM2.5 [277] https://taqm.epa.gov.tw/taqm/en/default.aspx
TaoBao [155] https://tianchi.aliyun.com/competition/entrance/
T-drive [244] https://www.kaggle.com/datasets/arashnic/tdriver
Tiingo [453] https://www.tiingo.com/
Traffic [177], [243], [118], [40], [287], [355], [186], [397], [138], [125], [424], [61], [271], [153], [451], [268] https://github.com/laiguokun/multivariate-time-series-data
UCI Beijing PM2.5 [505] https://archive.ics.uci.edu/ml/datasets/Beijing+PM2.5+Data
UCI human activity recognition [401], [456] https://archive.ics.uci.edu/dataset/240/human+activity+recognition+using+smartphones
UCR datasets [472] https://www.cs.ucr.edu/~eamonn/time_series_data_2018/
University of East Anglia (UEA) datasets [74], [113], [261], [140], [456] https://timeseriesclassification.com/
US historical climatology network (USHCN) [436], [491], [300] https://www.osti.gov/biblio/1389426
US stock market price [57] http://dx.doi.org/10.24963/ijcai.2019/810 (source paper)
Water distribution (WaDi) [396] https://itrust.sutd.edu.sg/itrust-labs_datasets/
WESAD dataset [337] https://ubi29.informatik.uni-siegen.de/usi/data_wesad.html
Wiki [451] https://www.kaggle.com/c/web-traffic-time-series-forecasting/data
Wikipedia network [232] http://snap.stanford.edu/jodie/wikipedia.csv
Windmill [327] https://pytorch-geometric-temporal.readthedocs.io/en/latest/notes/introduction.html#benchmark-datasets
Wind-speed [504] https://www.kaggle.com/datasets/fedesoriano/wind-speed-prediction-dataset
Wind generation [240] https://www.kaggle.com/datasets/sohier/30-years-of-european-wind-generation
@@
WISDM [100], [289] https://archive.ics.uci.edu/dataset/507/wisdm+smartphone+and+smartwatch+activity+and+biometrics+dataset
@@
The University of East Anglia (UEA) dataset includes 30 multivariate time series from many real-world applications, including heart records, electroencephalogram signals, motion data, audio spectra, written characters. Since a large portion of the dataset contains data not related to the focus of this review, the details of the results and models used will not be discussed here. Interested readers are referred to Refs. [74, 113, 261, 140, 456].